# Nonlinear Bayesian Filtering with Natural Gradient Gaussian Approximation

### v3

@@ Abstract
Index Terms:
State estimation, Bayesian filtering, Gaussian filter, natural gradient descent
@@ I Introduction
Notation: All vectors are considered as column vectors. The symbol $D_{\mathrm{KL}}(p\|q)=\mathbb{E}_{p}\left\{\log{\frac{p}{q}}\right\}$ denotes the Kullback-Leibler (KL) divergence between two probability distributions $p$ and $q$. Unless otherwise specified, $\|x\|$ refers to the $\ell_{2}$-norm of the vector $x$, defined as $\|x\|=\sqrt{x^{\top}x}$. The notation $\mathcal{N}(x;\mu,\Sigma)$ represents the Gaussian probability density function for the variable $x$ with mean $\mu$ and covariance matrix $\Sigma$. For simplicity, this may be abbreviated as $\mathcal{N}(\mu,\Sigma)$. Furthermore, $\mathbb{E}\{x\}$, $\mathbb{D}\{x\}$, and $\mathrm{Cov}(x,y)$ denote the expectation, variance, and covariance, respectively. The notation $A\leq B$ ($A<B$) for symmetric matrices $A$ and $B$ indicates that $B-A$ is positive (semi-)definite. The symbol $\otimes$ denotes the Kronecker product. The notation $\operatorname*{\mathrm{vec}}(A)$ refers to the vector obtained by stacking the columns of matrix $A$ into a single column vector. The trace of matrix $A$ is denoted as $\mathrm{Tr}(A)$. The notation $\mathbb{I}_{n\times n}$ represents the $n$-dimensional identity matrix.
@@ II Problem Statement
which is usually achieved through enabling approximation framework. In particular, this framework first approximates the transition or output probabilities with linear Gaussian models and then performs the well-known KF. Due to the closure property of Gaussians under linear transformations and the conjugate property of Gaussians under conditioning, the resulting prior and posterior distributions are naturally approximated as Gaussian. The linearization techniques in existing Gaussian filters can be categorized into two types: Taylor series expansion and stochastic linear regression. As summarized in Table I, the former uses the Jacobian matrix $g^{\prime}(\bar{x})$ to provide local affine approximations of system models at point $\bar{x}$, while the latter minimizes the expectation of the square of the residual $y-Ax-b$ to find optimal linear parameters.
@@
Taylor Series Expansion $g(x)=g(\bar{x})+g^{\prime}(\bar{x})(x-\bar{x})$Downarrow $A=g^{\prime}(\bar{x}),b=g(\bar{x})-g^{\prime}(\bar{x})\bar{x},\Lambda=\Sigma$ EKF[6], IEKF[8]
Statistical Linear Regression $\operatorname*{arg\,min}_{A,b}\mathbb{E}_{x}\left\{(y-Ax-b)^{\top}(y-Ax-b)\right\}$Downarrow $A=\bm{\mathrm{Cov}}(x,y)^{\top}\mathbb{D}(x),b=\mathbb{E}\left\{y\right\}-A\mathbb{E}\left\{x\right\},\Lambda=\mathbb{D}\left\{y\right\}-A\mathbb{D}\left\{x\right\}A^{\top}$ UKF [11], Gauss–Hermite KF [12], cubature KF [13], and PLF[14]
@@ III-A Optimization Viewpoint of Bayesian Filtering
This proposition shows that the prior distribution $p(x_{t}|y_{1:t-1})$ can be viewed as the solution to a variational problem that maximizes the expected logarithm of a candidate density $q(x_{t})$ over the joint distribution of the previous state and the transition probability. This reflects the idea that the prior is derived by considering all possible transitions from the previous state, and selecting the distribution that maximizes the expected log-density under these transitions. Similarly, the posterior distribution $p(x_{t}|y_{1:t})$ is the solution to a variational problem that minimizes a cost combining the expected negative log-likelihood of the measurement model, $-\log{p(y_{t}|{x}_{t})}$, with the KL divergence between the candidate density and the prior distribution $D_{\mathrm{KL}}(q(x_{t})||p(x_{t}|y_{1:t-1}))$. This captures the Bayesian update process, where the posterior distribution adjusts the prior distribution based on new measurements to balance prior knowledge with new information.
@@ III-B Structure of Stationary Points of $L(\hat{x}_{t},P_{t})$
For linear Gaussian systems $x_{t}=Ax_{t-1}+\xi_{t-1}$ where the process noise satisfies $\xi_{t-1}\sim\mathcal{N}(\xi_{t-1};0,Q)$, if the posterior distribution at time $t-1$ is given by $x_{t-1}\sim\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})$, the prediction step using (13) yields the following stationary points:
@@ V-B Stability Analysis
×$\times(y_{t}^{l}-g^{l}(\hat{x}_{t|t}+S_{t|t}z))\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
@@
×$\qquad\times\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
@@
×$\times(g^{l}(x)-g^{l}(\hat{x}_{t|t}+S_{t|t}z))\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
@@
×$\times\biggl{[}h_{j}^{kl}(\hat{x}_{t|t})+\frac{\partial h_{j}^{kl}}{\partial x^{q}}(x_{t}^{q}-\hat{x}_{t|t}^{q})+\psi_{j}^{kl}(x_{t}-\hat{x}_{t|t})\biggr{]}$
@@
×$\times\biggl{[}h_{j}^{kl}(\hat{x}_{t|t})+\frac{\partial h_{j}^{kl}}{\partial x^{q}}(\hat{x}_{t|t})e_{t|t}^{q}+\psi_{j}^{kl}(x_{t}-\hat{x}_{t|t})\biggr{]}$
@@
×$\times\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
@@
$0<\underline{p}I\leq\tilde{P}_{t|t-1}\leq\bar{p}I<\infty,\ \forall\ t\geq 0$
@@
×$\quad\quad\times\left[\left(F_{t-1}\tilde{P}_{t-1|t-1}F_{t-1}^{\top}+Q_{t}\right)^{-1}+D_{t}\right]$
×$\quad\quad\times(I+H_{t})^{-1}F_{t-1}e_{t-1|t-1}$
@@
×$\qquad\times\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\left(\Psi_{t}+\xi_{t}-G_{t}\zeta_{t}\right)$
@@
×$\times(I+H_{t})^{-1}F_{t-1}\leq\frac{1}{1+\lambda_{1}}\tilde{P}_{t-1|t-1}.$
@@ VI Robustifying NANO Filter with Gibbs Posterior
$=$ $\begin{cases}-\log{\mathcal{N}(y_{t};g(x_{t}),1)}&\text{if}\quad|y_{t}-g(x_{t})|\leq\delta,\\ -\log{\mathrm{Laplace}(y_{t};g(x_{t}),\frac{1}{\delta})}&\text{otherwise.}\end{cases}$
=$\stackrel{{\scriptstyle c}}{{=}}$$\begin{cases}\frac{1}{2}|y_{t}-g(x_{t})|^{2}&\text{if}\quad|y_{t}-g(x_{t})|\leq\delta,\\ \delta\left(|y_{t}-g(x_{t})|-\frac{1}{2}\delta\right)&\text{otherwise.}\end{cases}$
Here, the notation ``=$\stackrel{{\scriptstyle c}}{{=}}"$ represents that two expressions are equivalent up to an additive constant, and $\delta>0$ is the threshold variable. In fact, $\ell^{\mathrm{h}}(x_{t},y_{t})$ represents the famous Huber loss [38] used in robust regression, which is known to be less sensitive to outliers in data than the squared error loss. In practical applications, we utilize a smooth approximation of the Huber loss, known as the Pseudo-Huber loss, to improve optimization efficiency and ensure differentiability. The Pseudo-Huber loss is defined as:
@@ VIII Experiments
In this section, we validate the proposed NANO filter and its robust variants using the loss functions introduced in Section VI. We conduct simulations for both linear and nonlinear systems, along with a real-world experiment [^1] The code of this paper can be found in https://github.com/TianyiMehdi/NANO-filter. . All the simulations and experiment are evaluated by root mean square error (RMSE) which is defined by

### v4

@@ I Introduction
Notation: All vectors are considered as column vectors. The symbol $D_{\mathrm{KL}}(p\|q)=\mathbb{E}_{p}\left\{\log{\frac{p}{q}}\right\}$ denotes the Kullback-Leibler (KL) divergence between two probability distributions $p$ and $q$. Unless otherwise specified, $\|x\|$ refers to the $\ell_{2}$-norm of the vector $x$, defined as $\|x\|=\sqrt{x^{\top}x}$. The notation $\mathcal{N}(x;\mu,\Sigma)$ represents the Gaussian probability density function for the variable $x$ with mean $\mu$ and covariance matrix $\Sigma$. For simplicity, this may be abbreviated as $\mathcal{N}(\mu,\Sigma)$. Furthermore, $\mathbb{E}\{x\}$, $\mathbb{D}\{x\}$, and $\mathrm{Cov}(x,y)$ denote the expectation, variance, and covariance, respectively. The notation $A\leq B$ ($A<B$) for symmetric matrices $A$ and $B$ indicates that BAB-A is positive (semi-)definite. The symbol $\otimes$ denotes the Kronecker product. The notation $\operatorname*{\mathrm{vec}}(A)$ refers to the vector obtained by stacking the columns of matrix $A$ into a single column vector. The trace of matrix $A$ is denoted as $\mathrm{Tr}(A)$. The notation $\mathbb{I}_{n\times n}$ represents the $n$-dimensional identity matrix.
@@ II Problem Statement
which is usually achieved through enabling approximation framework. In particular, this framework first approximates the transition or output probabilities with linear Gaussian models and then performs the well-known KF. Due to the closure property of Gaussians under linear transformations and the conjugate property of Gaussians under conditioning, the resulting prior and posterior distributions are naturally approximated as Gaussian. The linearization techniques in existing Gaussian filters can be categorized into two types: Taylor series expansion and stochastic linear regression. As summarized in Table I, the former uses the Jacobian matrix $g^{\prime}(\bar{x})$ to provide local affine approximations of system models at point $\bar{x}$, while the latter minimizes the expectation of the square of the residual yAxby-Ax-b to find optimal linear parameters.
@@
Taylor Series Expansion $g(x)=g(\bar{x})+g^{\prime}(\bar{x})(x-\bar{x})$ $\Downarrow$ $A=g^{\prime}(\bar{x}),b=g(\bar{x})-g^{\prime}(\bar{x})\bar{x},\Lambda=\Sigma$ EKF[54], IEKF[22]
Statistical Linear Regression $\operatorname*{arg\,min}_{A,b}\mathbb{E}_{x}\left\{(y-Ax-b)^{\top}(y-Ax-b)\right\}$ $\Downarrow$ $A=\bm{\mathrm{Cov}}(x,y)^{\top}\mathbb{D}(x),b=\mathbb{E}\left\{y\right\}-A\mathbb{E}\left\{x\right\},\Lambda=\mathbb{D}\left\{y\right\}-A\mathbb{D}\left\{x\right\}A^{\top}$ UKF [32], Gauss–Hermite KF [4], cubature KF [5], and PLF[21]
@@ III-A Optimization Viewpoint of Bayesian Filtering
This proposition shows that the prior distribution $p(x_{t}|y_{1:t-1})$ can be viewed as the solution to a variational problem that maximizes the expected logarithm of a candidate density $q(x_{t})$ over the joint distribution of the previous state and the transition probability. This reflects the idea that the prior is derived by considering all possible transitions from the previous state, and selecting the distribution that maximizes the expected log-density under these transitions. Similarly, the posterior distribution $p(x_{t}|y_{1:t})$ is the solution to a variational problem that minimizes a cost combining the expected negative log-likelihood of the measurement model, $\log{p(y_{t}|{x}_{t})}$, with the KL divergence between the candidate density and the prior distribution $D_{\mathrm{KL}}(q(x_{t})||p(x_{t}|y_{1:t-1}))$. This captures the Bayesian update process, where the posterior distribution adjusts the prior distribution based on new measurements to balance prior knowledge with new information.
@@ III-B Structure of Stationary Points of $L(\hat{x}_{t},P_{t})$
For linear Gaussian systems $x_{t}=Ax_{t-1}+\xi_{t-1}$ where the process noise satisfies $\xi_{t-1}\sim\mathcal{N}(\xi_{t-1};0,Q)$, if the posterior distribution at time t1t-1 is given by $x_{t-1}\sim\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})$, the prediction step using (13) yields the following stationary points:
@@ V-B Stability Analysis
$\times(y_{t}^{l}-g^{l}(\hat{x}_{t|t}+S_{t|t}z))\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
@@
$\qquad\times\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
@@
$\times(g^{l}(x)-g^{l}(\hat{x}_{t|t}+S_{t|t}z))\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
@@
$\times\biggl[h_{j}^{kl}(\hat{x}_{t|t})+\frac{\partial h_{j}^{kl}}{\partial x^{q}}(x_{t}^{q}-\hat{x}_{t|t}^{q})+\psi_{j}^{kl}(x_{t}-\hat{x}_{t|t})\biggr]$
@@
$\times\biggl[h_{j}^{kl}(\hat{x}_{t|t})+\frac{\partial h_{j}^{kl}}{\partial x^{q}}(\hat{x}_{t|t})e_{t|t}^{q}+\psi_{j}^{kl}(x_{t}-\hat{x}_{t|t})\biggr]$
@@
$\times\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
@@
$0<\underline{p}I\leq\tilde{P}_{t|t-1}\leq\bar{p}I<\infty,\ \forall\ t\geq 0.$
@@
$\quad\quad\times\left[\left(F_{t-1}\tilde{P}_{t-1|t-1}F_{t-1}^{\top}+Q_{t}\right)^{-1}+D_{t}\right]$
$\quad\quad\times(I+H_{t})^{-1}F_{t-1}e_{t-1|t-1}$
@@
$\qquad\times\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\left(\Psi_{t}+\xi_{t}-G_{t}\zeta_{t}\right)$
@@
$\times(I+H_{t})^{-1}F_{t-1}\leq\frac{1}{1+\lambda_{1}}\tilde{P}_{t-1|t-1}.$
@@ VI Robustifying NANO Filter with Gibbs Posterior
[…]
Here, the notation $\stackrel{{\scriptstyle c}}{{=}}$" represents that two expressions are equivalent up to an additive constant, and $\delta>0$ is the threshold variable. In fact, $\ell^{\mathrm{h}}(x_{t},y_{t})$ represents the famous Huber loss [29] used in robust regression, which is known to be less sensitive to outliers in data than the squared error loss. In practical applications, we utilize a smooth approximation of the Huber loss, known as the Pseudo-Huber loss, to improve optimization efficiency and ensure differentiability. The Pseudo-Huber loss is defined as:
@@ VIII Experiments
In this section, we validate the proposed NANO filter and its robust variants using the loss functions introduced in Section VI. We conduct simulations for both linear and nonlinear systems, along with a real-world experiment . All the simulations and experiment are evaluated by root mean square error (RMSE) which is defined by